# On generalized Beauville decompositions

### v1

@@ Abstract.
On the other hand, we construct for any $g\geq 2$ a compactified Jacobian fibration of genus $g$ curves such that each curve is integral with at worst simple nodes and the (multiplicative) perverse filtration does not admit a multiplicative splitting. This shows that in general an extension of the Beauville decomposition cannot exist for compactified Jacobian fibrations even when the simplest singular point appears.
@@ 0. Introduction
Throughout, we work over the complex numbers ${\mathbb{C}}$. To motivate the discussion, we consider a principally polarized abelian scheme $\pi:A\to B$ of relative dimension $g$ so that $A$ is isomorphic to its dual $A^{\vee}=\mathrm{Pic}^{0}(A)$. The Fourier–Mukai transform
@@
associated with the normalized Poincaré line bundle $\mathcal{L}$ induces a Fourier transform
@@
The splitting $H^{*}(A,{\mathbb{Q}})=\oplus_{i=0}^{2g}H_{(i)}^{*}(A,{\mathbb{Q}})$ in Theorem 0.1, which is now referred to as the Beauville decomposition, is induced by the “multiplication by $N$” map $[N]:A\to A$; more precisely, the component $H_{(i)}^{*}(A,{\mathbb{Q}})$ is given by the eigenspace
@@
Furthermore, it has actually been shown in [7, 16] that the Beauville decomposition admits a motivic lifting.
The work of Arinkin [3, 4] extends the Fourier–Mukai transform (1) to compactified Jacobian fibrations associated with families of integral projective locally planar curves. Let $C\to B$ be such a family of curves over a nonsingular base $B$, and let $\pi:\overline{J}_{C}\to B$ be the associated compactified Jacobian fibration. [^1] Recall that the compactified Jacobian of an integral curve is the moduli space of rank $1$ degree 00 torsion free sheaves on that curve; it is integral and contains an open subset parameterizing line bundles when the singularities of the curve are planar. We further assume that the total space $\overline{J}_{C}$ is nonsingular. Analogously to the abelian scheme case, Arinkin’s Fourier–Mukai transform induces a (cohomological) Fourier transform
@@
When $C\to B$ arises as curves on a nonsingular surface $S$, we may view $\overline{J}_{C}$ as the moduli of certain stable torsion sheaves on $S$, and the Fourier transform $\mathfrak{F}$ governs tautological classes on $\overline{J}_{C}$; see [33, Section 5]. Therefore, a systematic study of Question 0.2 may lead to a better understanding of the “Perverse = Chern” phenomenon identifying perverse and Chern filtrations; this phenomenon was originally found for Hitchin systems via non-abelian Hodge theory and the $P=W$ conjecture [13, 32, 22] but was recently discovered in cases [25] beyond Hitchin moduli spaces.
@@
If we weaken the splitting to the filtration, then versions of Theorem 0.3(a, b) were proven to hold in [33] for any $\pi:\overline{J}_{C}\to B$ with $\overline{J}_{C}$ nonsingular. More precisely, it is shown that the perverse filtration is multiplicative and the Fourier transform is compatible with the perverse filtration; we refer to [33, Theorem 2.4] for the more precise statements. The point of Theorem 0.4 is that upgrading these properties from filtrations to decompositions seems to impose strong constraints on the family. For example, the proof of Theorem 0.3 relies heavily on recent developments in the study of derived categories of compact hyper-Kähler varieties. We expect that Question 0.2 has a positive answer when $\pi:\overline{J}_{C}\to B$ is Lagrangian. [^3] Here we do not require that $\overline{J}_{C}$ is compact. We refer to Section 1.5 for more discussions.
@@ 1.1. Fourier transforms
${\mathcal{P}}^{-1}:={\mathcal{H}}{\mathrm{om}}\left({\mathcal{P}},{\mathcal{O}}_{\overline{J}_{C}\times_{B}\overline{J}_{C}}\right)\otimes p_{2}^{*}\omega_{\pi}$
@@
where $T_{\overline{J}_{C}\times_{B}\overline{J}_{C}}$ is the virtual tangent bundle of the l.c.i. scheme $\overline{J}_{C}\times_{B}\overline{J}_{C}$. Since in general the relative product $\overline{J}_{C}\times_{B}\overline{J}_{C}$ is singular and the normalized Poincaré sheaf ${\mathcal{P}}$ of [4] does not admit a finite resolution by locally free sheaves, we use the tau-class $\tau(-)$ of [20, Chapter 18]; see also [33, Section 2.3] for a brief review.
@@
In particular, the Fourier transforms (3) on $H^{*}(\overline{J}_{C},{\mathbb{Q}})$ agree with the ones induced by the Fourier–Mukai kernels on the nonsingular absolute product $\overline{J}_{C}\times\overline{J}_{C}$ supported on the closed subset $\overline{J}_{C}\times_{B}\overline{J}_{C}$. We emphasize that the Fourier transforms are naturally relative correspondences over $B$. This is will be used to show Theorem 0.5(b) for the operators $f_{0},h_{0}$. We will discuss its motivic lifting in Section 3, where all the constructions are relative over $B$.
@@ 1.2. Decompositions
The first approach, as described in Section 0. Introduction, is to use the “multiplication by $N$” map $[N]:J_{C}\to J_{C}$ to decompose the total cohomology into eigenspaces. However, when singular fibers appear, it is difficult to extend the map $[N]$ over them.
The second approach, which was initiated in [7, 16], is to use the cycles (4) associated with the Fourier transforms to obtain projectors, which yields a (motivic) decomposition of the total cohomology. When there are singular fibers, this idea has been exploited in [33] intensively. It yields a number of consequences concerning the perverse filtration, which include the construction of a motivic decomposition splitting the perverse filtration analogously to the smooth case. However, as discussed in [33, Section 2.5.5], singular fibers force us to modify the correspondences constructed from $\mathfrak{F},\mathfrak{F}^{-1}$ in order to get projectors. Consequently, the resulting decomposition becomes mysterious and is therefore hard to analyze; in particular it is unclear whether it is multiplicative and how it interacts with the Fourier transform.
@@ 1.4. A proposal for constructing the $\mathfrak{sl}_{2}$-triple
Do $(e_{0},h_{0},f_{0})$ form an $\mathfrak{sl}_{2}$-triple with $f_{0}$ of degree $-2$?
Both the $\mathfrak{sl}_{2}$-triple and the degree $-2$ conditions concern a set of cohomology relations. If the answer to Question 1.4 is affirmative for $C\to B$, we get a decomposition from this $\mathfrak{sl}_{2}$-triple. Finally, we require that the decomposition obtained from the $\mathfrak{sl}_{2}$-triple $(e_{0},h_{0},f_{0})$ satisfies all the desired properties — it is Fourier-stable, splits the perverse filtration, and is multiplicative with respect to the cup-product.
@@ 2.1. Compact hyper-Kähler varieties
$[\mathrm{supp}({\mathcal{F}})]=L,\quad\chi({\mathcal{F}})=1-g,$
@@
By specialization, it suffices to solve Question 1.2 for a very general pair $(S,L)$. In this case, the moduli space $\overline{J}_{C}$ is of Picard rank $2$; the algebraic part of the cohomology is spanned by an ample class $A$ and the pullback of the hyperplane class $H$ of the base $B\simeq{\mathbb{P}}^{g}$.
We are looking for a solution to the following equation
$(A+\lambda H)^{g+1}=0,\quad\lambda\in{\mathbb{Q}}.$
Although this is a polynomial equation of degree $g$ (this is because $H^{g+1}=0$), the Beauville–Bogomolov–Fujiki (BBF) quadratic form $q(-)$ on $H^{2}(\overline{J}_{C},{\mathbb{Q}})$ [6] reduces it to a linear equation:
$q(A+\lambda H)=q(A)+2\lambda(A,H)=0.$
Here (,)(-,-)( - , - ) is the bilinear form associated with the BBF form, and we have used that $q(H)=0$. This solves $\lambda$ uniquely:
$\lambda=-\frac{q(A)}{2(A,H)}\in{\mathbb{Q}}.$
Now we consider arbitrary $(S,L)$ with $L$ an irreducible curve class. After changing the basis and specialization, we may assume that the algebraic part of the cohomology $H^{2}(\overline{J}_{C},{\mathbb{Q}})$ contains a relatively ample class $\Theta$ which, together with $H$, forms a rank $2$ hyperbolic ${\mathbb{Q}}$-lattice under the BBF form,
$q(\Theta)=q(H)=0,\quad(\Theta,H)=1.$
@@
The operator $f_{0}$ is of degree $-2$, and $(e_{0},h_{0},f_{0})$ form an $\mathfrak{sl}_{2}$-triple in $\mathrm{End}\left(H^{*}(\overline{J}_{C},{\mathbb{Q}})\right)$.
@@ 2.3. LLV algebras
with a quadratic form $\widetilde{q}(-)$. The form $\widetilde{q}(-)$ restricts to the BBF form on $H^{2}(X,{\mathbb{C}})$, and the two classes $\alpha,\beta$ are orthogonal to $H^{2}(X,{\mathbb{C}})$ satisfying
@@
We describe an interesting $\mathfrak{sl}_{2}$-triple in $\mathfrak{g}(X)$ which plays a crucial role [45]. The holomorphic symplectic form $\sigma$ on $X$ produces a class $\sigma\in H^{2}(X,{\mathbb{C}})$. It further induces an operator
@@ 2.4. Lie algebra calculations
They generate $\mathfrak{sl}_{2}$-triples $(e_{\eta_{i}},h,f_{\eta_{i}})$ in the LLV algebra $\mathfrak{g}(X)$ (this is essentially deduced in [45]; see e.g. [41, Lemma 2.5]).
@@
    $[e_{\sigma_{ij}},f_{\overline{\sigma}_{ij}}]=[e_{\overline{\sigma}_{ij}},f_{{\sigma}_{ij}}]=0.$
@@
Alternatively, we notice that when $\eta_{i},\eta_{j},\eta_{k}$ form the Kähler classes associated with three complex structures, the classes $\sigma_{ij},\overline{\sigma}_{ij}$ are given by the holomorphic symplectic form and its complex conjugate with respect to the complex structure corresponding to $\eta_{k}$. Therefore (a, b, c) follow from the fact that $\sigma,\overline{\sigma}$ generate an $\mathfrak{sl}_{2}\times\mathfrak{sl}_{2}\subset\mathfrak{g}(X)$ given by Example 2.2 and its complex conjugate; the general case can be reduced to this special case using hyper-Kähler rotations as in [45]. ∎
@@
(8) $[H,L]=2L,\quad[H,\Lambda]=-2\Lambda.$
@@
where we have used $K_{ij}=-K_{ji}$, $[K_{ij},K_{jk}]=2K_{ik}$ in calculating $H$. We further notice that
@@
from which we may calculate the commutators $[H,L]$, $[H,\Lambda]$. The commutation relations (8) then follow. ∎
@@ 2.6. Fourier transforms
By Verbitsky’s construction of the LLV algebra [45], any element $e_{\eta}$ associated with $\eta\in H^{2}(X,{\mathbb{C}})$ lies in the LLV algebra. Therefore, $e_{0}\in\mathfrak{g}(\overline{J}_{C})$. Alternatively, this can be seen from [41, Lemma 2.5], since any class $\eta\in H^{2}(\overline{J}_{C},{\mathbb{C}})$ can be written as the $\eta=\eta^{\prime}-\eta^{\prime\prime}$ with $q(\eta),q(\eta^{\prime})$ nonzero. By definition and Theorem 2.7, we have
@@
see [44, Theorems 4.8, 4.9]. Taelman’s construction relies on certain numerical assumptions which are satisfied by all known examples of compact hyper-Kähler varieties including $\overline{J}_{C}$.
@@
(12) $\alpha,\beta,\Theta,H\in\widetilde{H}(\overline{J}_{C},{\mathbb{C}})$
is completely described by the following formulas of Beckmann.
@@
We have
$\mathfrak{F}^{\widetilde{H}}(\alpha)=(-1)^{g}\left(\Theta-\frac{g+1}{2}\beta\right),\quad\mathfrak{F}^{\widetilde{H}}(\beta)=(-1)^{g-1}H,$
$\mathfrak{F}^{\widetilde{H}}(\Theta)=(-1)^{g-1}\alpha-\frac{g+1}{2}H,\quad\mathfrak{F}^{\widetilde{H}}(H)=(-1)^{g}\beta.$
@@
The action of (11) on the classes (12) is established in [11, Section 10.2]. [^4] Here we corrected some signs in [11]. The calculation uses the theory of extended Mukai vectors [11, 31], which further relies on the Rozansky–Witten theory governing the square root of the Todd class (see [11, Section 3]). For our purpose, it suffices to check that the class $-\frac{\lambda}{2}+\frac{g-1}{4}f$ in Beckmann’s calculation coincides with $-\Theta$; this class actually is uniquely characterized by that it and $H$ form a hyperbolic lattice when $(S,L)$ is very general. ∎
@@
Recall the class of the symplectic form $\sigma\in H^{2}(\overline{J}_{C},{\mathbb{C}})$. For any $\mu\in\mathrm{HT}^{2}(\overline{J}_{C})$, the associated element $e^{\prime}_{\mu}(\sigma)$ defines naturally an element in the extended Mukai lattice. This induces an embedding of vector spaces
@@
whose image is ${\mathbb{C}}\alpha\oplus H^{1,1}(\overline{J}_{C})\oplus{\mathbb{C}}\beta$; see [31, Equation (6.10)]. The identification
@@
We recall the expression of the $\mathfrak{sl}_{2}$-triple of Example 2.2 in terms of $\alpha$.
Lemma 2.10 ([44, Lemma 2.9]).
@@
$e^{\prime}_{\alpha}=f_{\sigma}.$
From now on, we also write the symplectic $\mathfrak{sl}_{2}$-triple as
$(e^{\prime}_{\alpha},h_{\mathrm{hol}},f^{\prime}_{\alpha}):=(f_{\sigma},h_{\mathrm{hol}},e_{\sigma})\subset\mathfrak{g}(\overline{J}_{C}).$
Next, since the action of the Fourier transform on the extended Mukai lattice is induced by a Hodge isometry, it preserves (13).
@@
There is a constant $c=1$ or $-1$, such that for any $\eta$ in (13) we have
$\mathfrak{F}^{\mathfrak{g}}(e^{\prime}_{\eta})=ce^{\prime}_{\mathfrak{F}^{\widetilde{H}}(\eta)}\in\mathfrak{g}(\overline{J}_{C}).$
@@
This follows from the construction of $\mathfrak{F}^{\widetilde{H}}$ and the argument of [31, Lemma 6.5]. ∎
@@
This operator relates the two types of $\mathfrak{sl}_{2}$-triples considered by Looijenga–Lunts–Verbitsky [30, 45] and Taelman [44] by conjugation. We collect some useful facts from [31, Section 6.1] and [46].
@@
Parts (a, b) are in [31, Section 6.1], and (c) is given by [46, Proposition 9.7]. ∎
@@
We will use cst to denote a constant which is not important for us. So Proposition 2.9 implies
(15) $\overline{\Theta}=(-1)^{g}\Theta+\mathrm{cst}\cdot\beta,\quad\mathfrak{F}^{\widetilde{H}}(\overline{\Theta})=-\alpha+\mathrm{cst}\cdot H.$
Since all the calculations in this section only concern the four classes $\alpha,\beta,\Theta,H$, we may use the conjugation of Lemma 2.12 to reduce them to the Lie algebra model of Section 2.4 with the four classes $\eta_{1},\eta_{2},\eta_{3},\eta_{4}$. More precisely, we may assume that after applying the conjugation

### v2

@@ Abstract.
On the other hand, we construct for any $g\geq 2$ a compactified Jacobian fibration of genus $g$ curves such that each curve is integral with at worst simple nodes and the (multiplicative) perverse filtration does not admit a multiplicative splitting. Our argument relies on the recently established universal double ramification cycle relations. This shows that in general an extension of the Beauville decomposition cannot exist for compactified Jacobian fibrations even when the simplest singular point appears.
@@ 0. Introduction
Throughout, we work over the complex numbers $\mathbb{C}$}. To motivate the discussion, we consider a principally polarized abelian scheme $\pi:A\to B$ of relative dimension $g$ over a nonsingular base $B$, so that $A$ is identified with its dual $A^{\vee}$. [^1] Since we are mostly concerned with (compactified) Jacobian fibrations, we assume principal polarization to simplify some of the statements. The Fourier–Mukai transform [36]
@@
associated with the normalized Poincaré line bundle $\mathcal{L}$ induces a (cohomological) Fourier transform [6]
@@
The splitting $H^{*}(A,{\mathbb{Q}})=\oplus_{i=0}^{2g}H_{(i)}^{*}(A,{\mathbb{Q}})$ in Theorem 0.1, which is now referred to as the Beauville decomposition, is induced by the “multiplication by NN” map $N]:A\to A$; more precisely, the component $H_{(i)}^{*}(A,{\mathbb{Q}})$ is given by the eigenspace
@@
Furthermore, it has actually been shown in [8, 17] that the Beauville decomposition admits a motivic lifting induced by algebraic cycles and cycle relations.
The work of Arinkin [3, 4] extends the Fourier–Mukai transform (1) to compactified Jacobian fibrations associated with families of integral projective locally planar curves. Let $C\to B$ be such a family of curves over a nonsingular base $B$, and let $\pi:\overline{J}_{C}\to B$ be the associated compactified Jacobian fibration. [^2] Recall that the compactified Jacobian of an integral curve is the moduli space of rank $1$ degree 0 torsion free sheaves on that curve; it is integral and contains an open subset parameterizing line bundles when the singularities of the curve are planar. We further assume that the total space $\overline{J}_{C}$ is nonsingular. Analogously to the abelian scheme case, Arinkin’s Fourier–Mukai transform induces a Fourier transform
@@
When $C\to B$ arises as curves on a nonsingular surface $S$, we may view $\overline{J}_{C}$ as the moduli of certain stable torsion sheaves on $S$, and the Fourier transform $\mathfrak{F}$ governs tautological classes on $\overline{J}_{C}$; see [34, Section 5]. Therefore, a systematic study of Question 0.2 may lead to a better understanding of the “Perverse = Chern” phenomenon identifying perverse and Chern filtrations; this phenomenon was originally found for Hitchin systems via non-abelian Hodge theory and the $P=W$ conjecture [14, 33, 23] but was recently discovered in cases [26] beyond Hitchin moduli spaces.
@@
If we weaken the splitting to the filtration, then versions of Theorem 0.3(a, b) were proven to hold in [34] for any $\pi:\overline{J}_{C}\to B$ with $\overline{J}_{C}$ nonsingular. More precisely, it is shown that the perverse filtration is multiplicative and the Fourier transform is compatible with the perverse filtration; we refer to [34, Theorem 2.4] for the more precise statements. The point of Theorem 0.4 is that upgrading these properties from filtrations to decompositions seems to impose strong constraints on the family. For example, the proof of Theorem 0.3 relies heavily on recent developments in the study of derived categories of compact hyper-Kähler varieties. We expect that Question 0.2 has a positive answer when $\pi:\overline{J}_{C}\to B$ is Lagrangian. [^4] Here we do not require that $\overline{J}_{C}$ is compact. We refer to Section 1.5 for more discussions.
@@ 1.1. Fourier transforms
${\mathcal{P}}^{-1}:={\mathcal{H}}{\mathrm{om}}\left({\mathcal{P}},{\mathcal{O}}_{\overline{J}_{C}\times_{B}\overline{J}_{C}}\right)\otimes p_{2}^{*}\omega_{\pi}[g]$
@@
where $T_{\overline{J}_{C}\times_{B}\overline{J}_{C}}$ is the virtual tangent bundle of the l.c.i. scheme $\overline{J}_{C}\times_{B}\overline{J}_{C}$. Since in general the relative product $\overline{J}_{C}\times_{B}\overline{J}_{C}$ is singular and the normalized Poincaré sheaf $\mathcal{P}$} of [4] does not admit a finite resolution by locally free sheaves, we use the tau-class $\tau(-)$ of [21, Chapter 18]; see also [34, Section 2.3] for a brief review.
@@
In particular, the Fourier transforms (3) on $H^{*}(\overline{J}_{C},{\mathbb{Q}})$ agree with the ones induced by the Fourier–Mukai kernels on the nonsingular absolute product $\overline{J}_{C}\times\overline{J}_{C}$ supported on the closed subset $\overline{J}_{C}\times_{B}\overline{J}_{C}$. We emphasize that the Fourier transforms are naturally relative correspondences over $B$. This will be used to show Theorem 0.5(b) for the operators $f_{0},h_{0}$. We will discuss its motivic lifting in Section 3, where all the constructions are relative over $B$.
@@ 1.2. Decompositions
The first approach, as described in Section 0. Introduction, is to use the “multiplication by NN” map $N]:J_{C}\to J_{C}$ to decompose the total cohomology into eigenspaces. However, when singular fibers appear, it is difficult to extend the map $[N]$ over them.
The second approach, which was initiated in [8, 17], is to use the cycles (4) associated with the Fourier transforms to obtain projectors, which yields a (motivic) decomposition of the total cohomology. When there are singular fibers, this idea has been exploited in [34] intensively. It yields a number of consequences concerning the perverse filtration, which include the construction of a motivic decomposition splitting the perverse filtration analogously to the smooth case. However, as discussed in [34, Section 2.5.5], singular fibers force us to modify the correspondences constructed from $\mathfrak{F},\mathfrak{F}^{-1}$ in order to get projectors. Consequently, the resulting decomposition becomes mysterious and is therefore hard to analyze; in particular it is unclear whether it is multiplicative and how it interacts with the Fourier transform.
@@ 1.4. A proposal for constructing the $\mathfrak{sl}_{2}$-triple
Do $(e_{0},h_{0},f_{0})$ form an $\mathfrak{sl}_{2}$-triple with $f_{0}$ of degree 2-2?
Both the $\mathfrak{sl}_{2}$-triple and the degree 2-2 conditions concern a set of cohomology relations. If the answer to Question 1.4 is affirmative for $C\to B$, we get a decomposition from this $\mathfrak{sl}_{2}$-triple. Finally, we require that the decomposition obtained from the $\mathfrak{sl}_{2}$-triple $(e_{0},h_{0},f_{0})$ satisfies all the desired properties — it is Fourier-stable, splits the perverse filtration, and is multiplicative with respect to the cup-product.
@@ 2.1. Compact hyper-Kähler varieties
$\mathrm{supp}({\mathcal{F}})]=L,\quad\chi({\mathcal{F}})=1-g,$
@@
By specialization, it suffices to solve Question 1.2 for a very general pair $(S,L)$. In this case, the moduli space $\overline{J}_{C}$ is of Picard rank $2$; the algebraic part of the cohomology $H^{2}(\overline{J}_{C},{\mathbb{Q}})$ is spanned by an ample class $\mathsf{A}$ and the pullback of the hyperplane class $\mathsf{H}$ of the base $B\simeq{\mathbb{P}}^{g}$.
We set $\Theta=\mathsf{A}+\lambda\mathsf{H}$ for some $\lambda\in{\mathbb{Q}}$, and we are looking for a solution to the following equation
$\Theta^{g+1}=(\mathsf{A}+\lambda\mathsf{H})^{g+1}=0.$
Although this is a polynomial equation of degree $g$ (this is because $\mathsf{H}^{g+1}=0$), the Beauville–Bogomolov–Fujiki (BBF) quadratic form q()q(-) on $H^{2}(\overline{J}_{C},{\mathbb{Q}})$ [7] reduces it to a linear equation:
$q(\mathsf{A}+\lambda\mathsf{H})=q(\mathsf{A})+2\lambda(\mathsf{A},\mathsf{H})=0.$
Here (,)(-,-) is the bilinear form associated with the BBF form, and we have used that $q(\mathsf{H})=0$. This solves $\lambda$ uniquely:
$\lambda=-\frac{q(\mathsf{A})}{2(\mathsf{A},\mathsf{H})}\in{\mathbb{Q}}.$
Now we consider arbitrary $(S,L)$ with $L$ an irreducible curve class. After changing the basis and specialization, we may assume that the algebraic part of $H^{2}(\overline{J}_{C},{\mathbb{Q}})$ contains a relatively ample class $\Theta$ which, together with $\mathsf{H}$, forms a rank $2$ hyperbolic $\mathbb{Q}$}-lattice under the BBF form,
$q(\Theta)=q(\mathsf{H})=0,\quad(\Theta,\mathsf{H})=1.$
@@
The operator $f_{0}$ is of degree 2-2, and $(e_{0},h_{0},f_{0})$ form an $\mathfrak{sl}_{2}$-triple in $\mathrm{End}\left(H^{*}(\overline{J}_{C},{\mathbb{Q}})\right)$.
@@ 2.3. LLV algebras
[…]
@@
We describe an interesting $\mathfrak{sl}_{2}$-triple in $\mathfrak{g}(X)$ which plays a crucial role in [47]. The holomorphic symplectic form $\sigma$ on $X$ produces a class $\sigma\in H^{2}(X,{\mathbb{C}})$. It further induces an operator
@@ 2.4. Lie algebra calculations
They generate $\mathfrak{sl}_{2}$-triples $(e_{\eta_{i}},h,f_{\eta_{i}})$ in the LLV algebra $\mathfrak{g}(X)$ (this is essentially deduced in [47]; see e.g. [43, Lemma 2.5]).
@@
    $e_{\sigma_{ij}},f_{\overline{\sigma}_{ij}}]=[e_{\overline{\sigma}_{ij}},f_{{\sigma}_{ij}}]=0.$
@@
Alternatively, we notice that when $\eta_{i},\eta_{j},\eta_{k}$ form the Kähler classes associated with three complex structures, the classes $\sigma_{ij},\overline{\sigma}_{ij}$ are given by the symplectic form $\sigma$ and its complex conjugate with respect to the complex structure corresponding to $\eta_{k}$. Therefore (a, b, c) follow from the fact that $\sigma,\overline{\sigma}$ generate an $\mathfrak{sl}_{2}\times\mathfrak{sl}_{2}\subset\mathfrak{g}(X)$ given by Example 2.2 and its complex conjugate; the general case can be reduced to this special case using hyper-Kähler rotations as in [47]. ∎
@@
(8) $H,L]=2L,\quad[H,\Lambda]=-2\Lambda.$
@@
where we have used $K_{ij}=-K_{ji}$, $K_{ij},K_{jk}]=2K_{ik}$ in calculating $H$. We further notice that
@@
from which we may calculate the commutators $[H,L]$, $H,\Lambda]$. The commutation relations (8) then follow. ∎
@@ 2.6. Fourier transforms
By Verbitsky’s construction of the LLV algebra [47], any element $e_{\eta}$ associated with $\eta\in H^{2}(X,{\mathbb{C}})$ lies in the LLV algebra. Therefore, $e_{0}\in\mathfrak{g}(\overline{J}_{C})$. Alternatively, this can be seen from [43, Lemma 2.5], since any class $\eta\in H^{2}(\overline{J}_{C},{\mathbb{C}})$ can be written as the $\eta=\eta^{\prime}-\eta^{\prime\prime}$ with $q(\eta),q(\eta^{\prime})$ nonzero. By definition and Theorem 2.7, we have
@@
see [46, Theorems 4.10, 4.11]. Taelman’s construction relies on certain numerical assumptions which are satisfied by all known examples of compact hyper-Kähler varieties including $\overline{J}_{C}$.
@@
(12) $\alpha,\beta,\Theta,\mathsf{H}\in\widetilde{H}(\overline{J}_{C},{\mathbb{C}})$
is described by the following formulas of Beckmann.
@@
There is a constant $c_{0}=1$ for $g$ odd, and $c_{0}=1$ or 1-1 for $g$ even, such that
$\mathfrak{F}^{\widetilde{H}}(\alpha)=-c_{0}\left(\Theta-\frac{g+1}{2}\beta\right),\quad\mathfrak{F}^{\widetilde{H}}(\beta)=c_{0}\mathsf{H},$
$\mathfrak{F}^{\widetilde{H}}(\Theta)=c_{0}\left(\alpha-\frac{g+1}{2}\mathsf{H}\right),\quad\mathfrak{F}^{\widetilde{H}}(\mathsf{H})=-c_{0}\beta.$
@@
The action of (11) on the classes (12) is established in [12, Section 10.2]. [^5] Here we corrected some signs in [12, Section 10.2] up to a global $\pm 1$; the method of [12] does not determine the precise sign for $g$ even. The calculation uses the theory of extended Mukai vectors [12, 32], which further relies on the Rozansky–Witten theory governing the square root of the Todd class (see [12, Section 3]). For our purpose, it suffices to check that the class $\frac{1}{2}\lambda-\frac{g-1}{4}f$ in Beckmann’s calculation coincides with $\Theta$; this class actually is uniquely characterized by that it and $\mathsf{H}$ form a hyperbolic lattice when $(S,L)$ is very general. ∎
We also remark that the class of the symplectic form $\sigma\in\widetilde{H}(\overline{J}_{C},{\mathbb{C}})$ is fixed up to a sign by the action of (11).
Proposition 2.10.
There is a constant $c_{1}=1$ for $g$ even, and $c_{1}=1$ or 1-1 for $g$ odd, such that
$\mathfrak{F}^{\widetilde{H}}(\sigma)=c_{1}\sigma.$
Proof.
We consider the composition $\Phi_{\mathcal{P}}\circ\Phi_{\mathcal{P}}$ and its induced action on the extended Mukai lattice, which equals $\mathfrak{F}^{\widetilde{H}}\circ\mathfrak{F}^{\widetilde{H}}$ by the functoriality of [46, Theorems 4.10 and 4.11]. Recall from [4] that
$\Phi_{\mathcal{P}}\circ\Phi_{\mathcal{P}}=\nu^{*}(-)\otimes\pi^{*}\omega_{B}[-g]$
where $\nu:\overline{J}_{C}\to\overline{J}_{C}$ is the antisymplectic involution induced by fiberwise duality. In particular, for $g$ odd we have
$\mathfrak{F}^{\widetilde{H}}\circ\mathfrak{F}^{\widetilde{H}}(\sigma)=\sigma.$
The same holds for $g$ even which is due to the extra orientation appearing in [46, Theorem 4.11]. In fact, by specialization we may assume $(S,L)$ very general so that the algebraic part of $H^{2}(\overline{J}_{C},{\mathbb{C}})$ is of rank $2$. The derived equivalence $\Phi_{\mathcal{P}}\circ\Phi_{\mathcal{P}}$ sends the structure sheaf $\mathcal{O}_{\overline{J}_{C}}$ to a line bundle shifted by an even number, which after [12, Lemma 4.2] implies
$\mathrm{det}(\mathfrak{F}^{\widetilde{H}}\circ\mathfrak{F}^{\widetilde{H}})=1.$
This forces $\mathfrak{F}^{\widetilde{H}}\circ\mathfrak{F}^{\widetilde{H}}$ to act as $1$ (and not 1-1) on the transcendental part of $\widetilde{H}(\overline{J}_{C},{\mathbb{C}})$ since the latter is of odd rank (=21=21).
For now we have $\mathfrak{F}^{\widetilde{H}}(\sigma)=\pm\sigma$. To see that the sign is positive for $g$ even, we repeat the argument above with $\Phi_{\mathcal{P}}$. Again by [12, Lemma 4.2] we have
$\mathrm{det}(\mathfrak{F}^{\widetilde{H}})=1$
since $\Phi_{\mathcal{P}}$ sends a sufficiently ample line bundle on $\overline{J}_{C}$ to a sheaf of positive rank. Comparing with the formulas of Proposition 2.9, we find again that $\mathfrak{F}^{\widetilde{H}}$ must act as $1$ on the transcendental part of $\widetilde{H}(\overline{J}_{C},{\mathbb{C}})$. ∎
@@
We consider the class of the symplectic form $\sigma\in H^{2}(\overline{J}_{C},{\mathbb{C}})$. For any $\mu\in\mathrm{HT}^{2}(\overline{J}_{C})$, the associated element $e^{\prime}_{\mu}(\sigma)$ defines naturally an element in the extended Mukai lattice. This induces an embedding of vector spaces
@@
whose image is ${\mathbb{C}}\alpha\oplus H^{1,1}(\overline{J}_{C})\oplus{\mathbb{C}}\beta$; see [32, Equation (6.10)]. The identification
@@
The $\mathfrak{sl}_{2}$-triple of Example 2.2 can now be expressed in terms of $\alpha$.
Lemma 2.11 ([46, Lemma 2.9]).
@@
$e^{\prime}_{\alpha}=f_{\sigma}\in\mathfrak{g}(\overline{J}_{C}).$
We thus obtain an $\mathfrak{sl}_{2}$-triple
(14) $(e^{\prime}_{\alpha},-h_{\mathrm{hol}},f^{\prime}_{\alpha})=(f_{\sigma},-h_{\mathrm{hol}},-e_{\sigma})\subset\mathfrak{g}(\overline{J}_{C}).$
Next, since the action of the Fourier transform (11) on the extended Mukai lattice is induced by a Hodge isometry, it preserves (13). We have the following comparison.
@@
Let $c_{1}$ be the constant in Proposition 2.10. For any $\eta$ in (13), we have
$\mathfrak{F}^{\mathfrak{g}}(e^{\prime}_{\eta})=c_{1}e^{\prime}_{\mathfrak{F}^{\widetilde{H}}(\eta)}\in\mathfrak{g}(\overline{J}_{C}).$
@@
This follows from [32, Lemma 6.5] together with Proposition 2.10. ∎
@@
This operator relates the two types of $\mathfrak{sl}_{2}$-triples considered by Looijenga–Lunts–Verbitsky [31, 47] and Taelman [46] by conjugation. We collect some useful facts from [32, Section 6.1] and [48].
@@
Parts (a, b) are in [32, Section 6.1], and (c) is given by [48, Proposition 9.7]. ∎
@@
We will use $\mathrm{cst}$” to denote a constant which is not important for us. So Proposition 2.9 reads
(16) $\overline{\Theta}=-c_{0}\Theta+\mathrm{cst}\cdot\beta,\quad\mathfrak{F}^{\widetilde{H}}(\overline{\Theta})=-\alpha+\mathrm{cst}\cdot\mathsf{H}.$
Since all the calculations in this section only concern the four classes $\alpha,\beta,\Theta,\mathsf{H}$, we may use the conjugation of Lemma 2.13 to reduce them to the Lie algebra model of Section 2.4 with the four classes $\eta_{1},\eta_{2},\eta_{3},\eta_{4}$. More precisely, we may assume that after applying the conjugation